An electron in an oscilloscope tube is traveling horizontally with a speed of 4.0
Answer #1

a = -1.602*10^-19 * 1700/9.11*10^-31 = -298.946 * 10^12 m/sec^2

v^2-u^2 = 2as

v^2 - (4*10^6)^2 = 2*298.946*10^12 * 0.021

v^2 = 1.2555732*10^13 + 1.6 * 10^13

v = 5.343756 * 10^6 m/s